OpenStudy (anonymous):

Find the local max and min values of f. \[f(x)=12-2\left| x+3 \right|\]

OpenStudy (michele_laino):

we have to explicit the absolute value

OpenStudy (michele_laino):

hint: \[\begin{gathered} \left| {x + 3} \right| = x + 3,\quad x \geqslant - 3 \hfill \\ \hfill \\ \left| {x + 3} \right| = - \left( {x + 3} \right),\quad x < - 3 \hfill \\ \end{gathered} \]

OpenStudy (jhannybean):



OpenStudy (anonymous):

How do I know which one to use? @Michele_Laino

OpenStudy (michele_laino):

both methods are good. nevertheless if you want to rewrite the equation of your function, you have to substitute the absolute value, as I have indicated above

OpenStudy (michele_laino):

for example, if \[x \geqslant - 3\] then your function can be rewritten as below: \[12 - 2\left( {x + 3} \right) = 12 - 2x - 6 = 6 - 2x\] which is represented by the right branch of the drawing of @Jhannybean whereas, if: \[x < - 3\] then your function can be rewritten as below: \[12 - 2\left[ { - \left( {x + 3} \right)} \right] = 12 + 2x + 6 = 18 + 2x\] which is represented by the left branch of the drawing of @Jhannybean

OpenStudy (anonymous):

oh okay, so then I take the derivatative of one of them? @Michele_Laino

OpenStudy (irishboy123):

mmm look at the graph do you need to do anything else to find max and mins?

OpenStudy (michele_laino):

It is suffice to keep in mind the graph of @Jhannybean

OpenStudy (michele_laino):

what is the maximum value?

OpenStudy (anonymous):

when f(-3)=12 @Michele_Laino

OpenStudy (anonymous):

from the graph... But would you get the -3 from?

OpenStudy (anonymous):

algebraically?

OpenStudy (michele_laino):

ok! That's right!